OpenStudy (anonymous):

Find the area of a rectangle with base of 12 inches and height of 8 inches

OpenStudy (llort):

l*w=a 12*8=your answer

OpenStudy (dumbcow):

12*8 = 96

OpenStudy (llort):

@everyone else: give a man a fish, he will eat for a day. teach a man to fish and he will eat for a lifetime.

OpenStudy (anonymous):

didnt know a triangle has a height

jhonyy9 (jhonyy9):

12 multiplied with 8 = area of rectangle

jhonyy9 (jhonyy9):

12x8=96